# A pure dual approach for hedging Bermudan options

### v1

@@ Introduction
This paper proposes a new algorithm to compute the hedging portfolio of a Bermudan option and its corresponding price. American and Bermudan options have been widely studied in the literature. The algorithm proposed by Longstaff and Schwartz (2001) is probably the most widely used in the financial industry for pricing these options by Monte Carlo. It consists first in approximating the optimal stopping rule and then in using it to evaluate the Bermudan claim. It is a pure primal algorithm that gives the value from the option buyer point of view. However, in the theory of Black, Scholes and Merton, the price of an option corresponds to the initial value of a self-financing portfolio which exactly hedges the option. In the case of Bermudan options, it is not clear for an option seller how to construct a hedging portfolio from the single Longstaff and Schwartz price.
@@
To come up with this issue, Rogers (2002) and Haugh and Kogan (2004) have introduced a dual formula that directly puts in evidence the hedging portfolio by the mean of a martingale. Rogers (2002) derived from this formula an algorithm which provides a hedging portfolio, but the algorithm is however limited to some specific payoff functions (see the discussion in Subsection 4.1 for an illustration). Haugh and Kogan (2004) and later on Andersen and Broadie (2004) used this dual formula to produce an upper bound of the Bermudan option price, while the Longstaff Schwartz algorithm provides a lower bound. To be more precise, Andersen and Broadie (2004) used the Longstaff Schwartz algorithm to construct a suitable martingale from the approximated optimal stopping time. This martingale produces their upper bound price. However, it is again not obvious to construct the hedging portfolio that corresponds to this martingale. This issue was pointed out by Rogers (2010) who advocated a purely dual approach, i.e. an algorithm that is only based on the dual formula. This is precisely the goal of our paper: we develop a purely dual algorithm that directly calculates the hedging portfolio of a Bermudan option.
@@
To obtain our algorithm, we derive two key results from the dual formula (Rogers, 2002, Theorem 2.1). First, we rewrite the expectation as a telescopic sum, which enables us to decompose the minimisation problem into a sequence of minimisation problems on the martingale increments. Second, we introduce a strict convexification of these problems, which gives the uniqueness of the solution and makes the resolution of the problem easily tractable. These two steps are explained in Section 1. The remainder of the paper is organized as follows. Section 2 analyses the approximation error of our purely dual algorithm and shows the convergence of the related Monte Carlo estimators. Section 3 presents the implementation of our algorithm in a financial framework and make precise the martingales that are used in the hedging strategy. Last, Section 4 gives a detailed presentation of our numerical results for the hedging and valuation of different Bermudan options on one or several assets. We show the relevance of the algorithm, reporting the P&L of discrete hedging portfolios calculated with it. Namely, we present hedging portfolios with different rebalancing frequencies that are either constructed with the underlying assets only or with adding vanilla European options. Of course, a hedging portfolio with a higher rebalancing frequency and with more financial instruments can in principle do better. However, this requires higher computational efforts and also in practice higher operational costs. Thus, it is interesting to assess the relevance of increasing the rebalancing frequency or of adding financial instruments in the hedging portfolio. We observe on our examples that there is no absolute answer. Adding European options in the hedging portfolio improves the hedge in most of our cases but one, while increasing the rebalancing frequency is relevant up to a certain value, depending on the option considered. Thus, beyond giving an upper bound price of the Bermudan option and its associated hedging portfolio, our algorithm is useful
@@ 1 Framework and main result
see (Lamberton and Lapeyre, 1996, Section 2.2) for instance. The process $U$ is a $L^{p}$ supermartingale, $\tau^{\star}:=\inf\{n\in\{0,\dots,N\}:U_{n}=Z_{n}\}$ is an optimal stopping time for (1), and $(U_{n\wedge\tau^{\star}}){0\leq n\leq N}$ is a martingale.
@@ 1.1 Excess reward representation of the optimal stopping problem
In the context of mathematical finance, (9) can be seen as the expansion of the Bermudan option price as a sum of the corresponding European option price ${\mathbb{E}}[Z_{N}]$ and of correcting terms ${\mathbb{E}}\left[\left(Z_{n}+\Delta M_{n+1}-\max_{n+1\leq j\leq N}\left\{Z_{j}-\sum_{i=n+2}^{j}\Delta M_{i}\right\}\right)_{+}\right]$ representing, for $M=M^{\star}$, the value of having the right to exercise the option at time $n\in\{0,\dots,N-1\}$.
@@ 1.2 Strict convexification and algorithm
$\inf_{\Delta M_{n+1}\in{\mathcal{H}}^{p}_{n+1}}{\mathbb{E}}\left[\left(Z_{n}+\Delta M_{n+1}-\max_{n+1\leq j\leq N}\left\{Z_{j}-\sum_{i=n+2}^{j}\Delta M_{i}\right\}\right)_{+}\right],$
@@
We can now describe formally our algorithm:
@@ 2.2 Convergence of the Monte Carlo estimators
Now, we assume that we have $Q$ independent paths of the underlying process $Z_{n}^{q}$ and of the martingale increments $\Delta X^{q}_{n}$, for $1\leq n\leq N$ and $1\leq q\leq Q$. Here, and through the paper, the index $q$ will denote the $q-$th sample of the Monte Carlo algorithm. Then, we minimize, for $n=N-1$ to $n=0$,
$\alpha^{Q}_{n+1}=\arg\min_{\alpha\in{\mathbb{R}}^{L}}\frac{1}{Q}\sum_{q=1}^{Q}\left(Z^{q}_{n}+\alpha\cdot\Delta X^{q}_{n+1}-\max_{n+1\leq j\leq N}\left\{Z^{q}_{j}-\sum_{i=n+2}^{j}\alpha^{Q}_{i}\cdot\Delta X^{q}_{i}\right\}\right)^{2}.$
Again, we write $\alpha\cdot\Delta X^{q}_{n+1}=\sum_{\ell=1}^{L}\alpha_{\ell}\Delta X^{q}_{n+1,\ell}$. If the matrix $\frac{1}{Q}\sum_{q=1}^{Q}\Delta X^{q}_{n+1}(\Delta X^{q}_{n+1})^{T}$ is invertible, then $\alpha^{Q}_{n+1}$ is unique and given by
@@ 4.1 One dimensional options
First, we comment the results with the stock only portfolio. As we may expect, the rebalancing frequency plays a crucial role. The higher is the frequency, the closer is the estimated dual price to the primal price given by the Longstaff and Schwartz algorithm. We observe on this example the effect of overfitting: for $\bar{N}=10$ and $P=100$, the total number of parameters to estimate is $N\times\bar{N}\times P=10000$, and $Q=100000$ is not sufficient for a suitable estimation, as shown by the difference between $U_{0}^{Q}$ and $\hat{U}_{0}^{Q}$. Increasing $Q$ to $500000$ reduces this gap. With $\bar{N}=20$, $P=50$ and $Q=2\times 10^{6}$, we get a price which is $0.06$ above the Longstaff-Schwartz price, which is satisfactory: the same error magnitude appears also when hedging European options with discrete time strategies. We now discuss the P&L obtained in Figure 1. We observe the importance of the rebalancing frequency for a stock only portfolio, and get a variance of 2.73 when $\bar{N}=5$ (Figure 1(a)) and $1.05$ when $\bar{N}=10$ (Figure 1(b)).
@@
Figure 1: P&L histograms of the hedging strategy for the Bermudan Put option of Table 1 for the stock only strategy (top) and the strategy using extra European options (bottom).
When adding the European put option in the hedging portfolio, our algorithm is very close to the primal price with already $\bar{N}=1$ and $P=1$ (note that considering $P=50$ does not help much). This is also confirmed in Figure 1(c) where the P&L has a variance equal to 0.13: most of the risk is hedged by the European options. This was already expected from the conclusions of Rogers (2002). More precisely, let us consider the following martingale increments between two exercising dates
@@
Thus, in the put example, scaling the martingale associated with the European claim gives already a nice approximation of $M^{\star}$, better than the one obtained with a discrete hedging strategy using only the asset. From a theoretical point of view, adding European options in the hedging portfolio amounts to allow some particular continuous time trading. For the put option, this is very efficient and if we do not include them, our algorithm essentially finds an optimal discrete approximation of this continuous time strategy.
@@
For the same Bermudan Butterfly option, Table 2 shows the result of our algorithm when using a hedging portfolio using only the asset (${\mathcal{A}}^{1}_{t}=\tilde{S}_{t}$) or using the asset and a European call option (${\mathcal{A}}^{1}_{t}=\tilde{S}_{t}$ and ${\mathcal{A}}^{2}_{t}=e^{-rT}{\mathbb{E}}[(S_{T}-\frac{K_{1}+K_{2}}{2})_{+}|{\mathcal{F}}_{t}]$ in (21)). We see again a clear advantage to include the European option in the hedging strategy. With $\bar{N}=10$ and $P=100$, we find a dual price equal to $5.82$. With $\bar{N}=20$ and $P=50$, we find a dual price equal to $5.74$, which is very close to the Longstaff-Schwartz price: doubling the rebalancing frequency allows to propose a smaller price (0.08 less). Thus, our algorithm does not only give a hedging strategy, it can also be used by practitioners to find a good trade-off between the price and the cost of extra monitoring.
To go beyond the prices, we have drawn in Figure 2 the P&L of the hedging strategy obtained with our algorithm. We see that the use of European options in the hedging strategy reduces the variance of the P&L: 1.13 versus 0.48. Thus, including European options does not only improve the selling price but it gives a lower variance of the P&L distribution.
Refer to captionRefer to caption

### v2

@@ Introduction
This paper proposes a new algorithm to compute the hedging portfolio of a Bermudan option and its corresponding price. American and Bermudan options have been widely studied in the literature. The algorithm proposed by Longstaff and Schwartz (2001) is probably the most widely used in the financial industry for pricing these options by Monte Carlo. It consists first in approximating the optimal stopping rule and then in using it to evaluate the Bermudan claim. It is a pure primal algorithm that gives the value from the option buyer point of view. Many different algorithms have been proposed for the valuation of American or Bermudan claims by Monte-Carlo. Here, we mention in particular Haugh and Kogan (2004); Andersen and Broadie (2004); Belomestny et al. (2009); Desai et al. (2012) who aimed at giving both upper and lower bounds for the Bermudan price. However, in the theory of Black, Scholes and Merton, the price of an option corresponds to the initial value of a self-financing portfolio which exactly hedges the option. In the case of Bermudan options, it is not clear for an option seller how to construct a practical hedging portfolio from the single Longstaff and Schwartz price or, more generally, from the prices obtained by the different methods mentioned above.
@@
Several papers such as Wang and Caflisch (2010); Belomestny et al. (2010) have proposed Monte-Carlo methods to compute the delta and other sensitivities of Bermudan options that can be used then for hedging. In a different direction, Rogers (2002) and Haugh and Kogan (2004) have introduced a dual formula that directly puts in evidence the hedging portfolio by the mean of a martingale. Rogers (2002) derived from this formula an algorithm which provides a hedging portfolio, but the algorithm is however limited to some specific payoff functions (see the discussion in Subsection 4.1 for an illustration). Desai et al. (2012) have proposed a more generic Monte-Carlo approach with a finite vector subspace of approximating martingales, so that the dual formula boils down to a Linear Program (LP). Even though LP can be solved very efficiently, there is a high number of constraints that restricts in practice the number of samples and/or the dimension of the martingale approximating space. However in the literature, the dual formula has mostly been used to produce an upper bound of the Bermudan option price as initiated by Haugh and Kogan (2004) and later on Andersen and Broadie (2004), while the Longstaff Schwartz algorithm provides a lower bound. To be more precise, Andersen and Broadie (2004) used the Longstaff Schwartz algorithm to construct a suitable martingale from the approximated optimal stopping time. This martingale produces their upper bound price. However, it is again not obvious how to construct a practical hedging portfolio that corresponds to this martingale by using tradable financial instruments. This issue was pointed out by Rogers (2010) who advocated a purely dual approach, i.e. an algorithm that is only based on the dual formula. This is precisely the goal of our paper: we develop a purely dual algorithm that directly calculates the hedging portfolio of a Bermudan option that can be built using only tradable financial instruments.
@@
To obtain our algorithm, we derive two key results from the dual formula (Rogers, 2002, Theorem 2.1). First, we rewrite the expectation as a telescopic sum, which enables us to decompose the minimisation problem into a sequence of minimisation problems on the martingale increments. This decomposition is a key difference with the LP algorithm of Desai et al. (2012) which directly solves the global minimisation problem. Second, we introduce a strict convexification of these problems, which gives the uniqueness of the solution and makes the resolution of the problem easily tractable. These two steps are explained in Section 1. The remainder of the paper is organized as follows. Section 2 analyses the approximation error of our purely dual algorithm and shows the convergence of the related Monte Carlo estimators. Section 3 presents the implementation of our algorithm in a financial framework and make precise the martingales that are used in the hedging strategy. Last, Section 4 gives a detailed presentation of our numerical results for the hedging and valuation of different Bermudan options on one or several assets. We show the relevance of the algorithm, comparing it with some existing delta hedging methods, and reporting the P&L of discrete hedging portfolios calculated with it. Namely, we present hedging portfolios with different rebalancing frequencies that are either constructed with the underlying assets only or with adding vanilla European options. Of course, a hedging portfolio with a higher rebalancing frequency and with more financial instruments can in principle do better. However, this requires higher computational efforts and also in practice higher operational costs. Thus, it is interesting to assess the relevance of increasing the rebalancing frequency or of adding financial instruments in the hedging portfolio. We observe on our examples that there is no absolute answer. Adding European options in the hedging portfolio improves the hedge in most of our cases but one, while increasing the rebalancing frequency is relevant up to a certain value, depending on the option considered. Thus, beyond giving an upper bound price of the Bermudan option and its associated hedging portfolio, our algorithm is useful
@@ 1 Framework and main result
see (Lamberton and Lapeyre, 1996, Section 2.2) for instance. The process $U$ is a $L^{p}$ supermartingale, $\tau^{\star}:=\inf\{n\in\{0,\dots,N\}:U_{n}=Z_{n}\}$ is an optimal stopping time for (1), and $(U_{n\wedge\tau^{\star}})_{0\leq n\leq N}$ is a martingale.
@@ 1.1 Excess reward representation of the optimal stopping problem
Therefore, we can rewrite the global minimisation as follows
$U_{0}$ $=\inf_{M\in{\mathbb{H}}^{p}}{\mathbb{E}}\left[\max_{0\leq j\leq N}\{Z_{j}-(M_{j}-M_{0})\}\right]$
$={\mathbb{E}}[Z_{N}]+\inf_{\Delta M_{1}\in{\mathcal{H}}_{1}^{p},\dots,\Delta M_{N}\in{\mathcal{H}}_{N}^{p}}\sum_{n=0}^{N-1}{\mathbb{E}}\left[\left(Z_{n}+\Delta M_{n+1}-\max_{n+1\leq j\leq N}\left\{Z_{j}-\sum_{i=n+2}^{j}\Delta M_{i}\right\}\right)_{+}\right].$ (10)
@@
In the context of mathematical finance, (9) can be seen as the expansion of the Bermudan option price as a sum of the corresponding European option price ${\mathbb{E}}[Z_{N}]$ and of correcting terms ${\mathbb{E}}\left[\left(Z_{n}+\Delta M_{n+1}-\max_{n+1\leq j\leq N}\left\{Z_{j}-\sum_{i=n+2}^{j}\Delta M_{i}\right\}\right)_{+}\right]$ representing, for $M=M^{\star}$, the value of having the right to exercise the option at time $n\in\{0,\dots,N-1\}$. From (7) and (6), we get indeed for $M=M^{\star}$ that
$\left(Z_{n}+\Delta M_{n+1}-\max_{n+1\leq j\leq N}\left\{Z_{j}-\sum_{i=n+2}^{j}\Delta M_{i}\right\}\right)_{+}=(Z_{n}-{\mathbb{E}}[U_{n+1}|\mathcal{F}_{n}])_{+},$
which measures the extra reward gained by exercising rather than continuing. The identity
$U_{0}={\mathbb{E}}[Z_{N}]+\sum_{n=0}^{N-1}{\mathbb{E}}[(Z_{n}-{\mathbb{E}}[U_{n+1}|\mathcal{F}_{n}])_{+}]$
can also be directly derived from the dynamic programming equation (2) and has been intensively used by Belomestny et al. (2009) to compute Bermudan price bounds. In contrast, (9) is derived from the dual formula and holds for any martingale $M$. This will be essential to decompose the minimisation problem (10) across the time.
@@ 1.2 Strict convexification and algorithm
$\inf_{\Delta M_{n+1}\in{\mathcal{H}}^{p}_{n+1}}{\mathbb{E}}\left[\left(Z_{n}+\Delta M_{n+1}-\max_{n+1\leq j\leq N}\left\{Z_{j}-\sum_{i=n+2}^{j}\Delta M_{i}\right\}\right)_{+}\right],$ (11)
@@
Remark 1.3.
Let us emphasize that there are many solutions of the backward minimization problem (11). Indeed, using (Schoenmakers et al., 2013, Theorem 4.1), we show below in Proposition 1.6 this result. To circumvent the non-uniqueness issue, (Schoenmakers et al., 2013, Theorem 5.1) proposes to add a conditional variance penalization in the minimization problem, that is then no longer convex. Here, we rather use a convexification technique that makes the minimization problem strictly convex.
@@
Proposition 1.6.
Let $M^{\star}$ and $A^{\star}$ be defined as in (6). Let $(\zeta_{n})_{1\leq n\leq N}$ be any sequence of adapted random variables such that ${\mathbb{E}}[\zeta_{n}|{\mathcal{F}}_{n-1}]=1$ for $n\geq 1$ and such that ${\mathbb{P}}(\zeta_{n}\geq 0|{\mathcal{F}}_{n-1})=1$. Then, the martingale
$M_{n}=M^{\star}_{n}-A^{\star}_{n}+\sum_{j=1}^{n}(A^{\star}_{n}-A^{\star}_{n-1})\zeta_{n},\ n\geq 1,$
solves the backward minimisation problem (11).
Proof.
Let us first recall that an integrable random variable $\xi$ satisfies ${\mathbb{E}}[\xi_{+}]={\mathbb{E}}[\xi]_{+}$ if ${\mathbb{P}}(\xi\geq 0)=1$ or $\xi$ is constant. Therefore, the equality in (14) with $\varphi(x)=x_{+}$ holds if
[…]
By (Schoenmakers et al., 2013, Theorem 4.1), we have $U_{n}=\max_{n\leq j\leq N}(Z_{j}-M_{j}+M_{n})$ for $0\leq n\leq N$, and thus
$Z_{n}+\Delta M_{n+1}-\max_{n+1\leq j\leq N}\left\{Z_{j}-\sum_{i=n+2}^{j}\Delta M_{i}\right\}$ $=Z_{n}+\Delta M_{n+1}-U_{n+1}$
$=Z_{n}-{\mathbb{E}}[U_{n+1}|{\mathcal{F}}_{n}]+\Delta A_{n+1}^{\star}(\zeta_{n+1}-1).$
If $\Delta A_{n+1}^{\star}>0$, we have $U_{n}=Z_{n}$. Since $U_{n}-{\mathbb{E}}[U_{n+1}|{\mathcal{F}}_{n}]=\Delta A_{n+1}^{\star}$, we get $Z_{n}-{\mathbb{E}}[U_{n+1}|{\mathcal{F}}_{n}]+\Delta A_{n+1}^{\star}(\zeta_{n+1}-1)=\Delta A_{n+1}^{\star}\zeta_{n+1}.$ This leads to
$Z_{n}-{\mathbb{E}}[U_{n+1}|{\mathcal{F}}_{n}]+\Delta A_{n+1}^{\star}(\zeta_{n+1}-1)=\mathbf{1}_{\Delta A_{n+1}^{\star}=0}\left(Z_{n}-{\mathbb{E}}[U_{n+1}|{\mathcal{F}}_{n}]\right)+\mathbf{1}_{\Delta A_{n+1}^{\star}>0}\Delta A_{n+1}^{\star}\zeta_{n+1}.$
Since $\Delta A_{n+1}^{\star}$ is ${\mathcal{F}}_{n}$-measurable, we get (16), which shows that $M$ solves (11). $\blacksquare$
Remark 1.7.
Proposition 1.6 shows that there are, in general, many solutions to the sequence of the backward minimisation problems (11). Then, it is not clear that a solution of this sequence solves the global minimisation problem (10). This shows a double interest of the strict convexification: it ensures first uniqueness of the solution of each problem (13), which gives then precisely the unique solution of the global minimisation problem (10).
Description of the algorithm. We now describe formally our algorithm:
@@ 2.2 Convergence of the Monte Carlo estimators
Now, we assume that we have $Q$ independent paths of the underlying process $Z_{n}^{q}$ and of the martingale increments $\Delta X^{q}_{n}$, for $1\leq n\leq N$ and $1\leq q\leq Q$. Here, and through the paper, the index $q$ will denote the $q-$th sample of the Monte Carlo algorithm. If the matrix $\frac{1}{Q}\sum_{q=1}^{Q}\Delta X^{q}_{n+1}(\Delta X^{q}_{n+1})^{T}$ is invertible, then we define $\alpha^{Q}_{n+1}$ as the Monte Carlo estimator of (20) by
@@ 3 Implementation in a financial framework
Remark 3.1.
[Computational complexity] In practice, we consider a Monte-Carlo estimator of (26) with $Q$ samples and we have to solve $N\times\bar{N}$ linear systems of dimension $\bar{P}\times\bar{d}$. Thus, the overall complexity is of order $O(QN\bar{N}(\bar{P}\bar{d})^{3})$, which can even be $O(QN\bar{N}\bar{P}\bar{d}^{3})$ when choosing a local regression basis. In contrast, the global minimisation proposed in Desai et al. (2012) requires to solve a Linear Program with $Q+N\times\bar{N}\times\bar{P}\times\bar{d}$ variables and $Q\times(N+1)$ constraints and is therefore far most computationally demanding, see e.g. Cohen et al. (2021) for recent results on the LP complexity.
@@ 4.1 One dimensional options
(d) $\bar{N}=5$, $P=50$, $Q=10^{5}$, delta hedging (CRR).
(e) $\bar{N}=5$, $P=50$, $Q=10^{5}$, delta hedging (Wang Caflisch).
Figure 1: P&L histograms of the hedging strategy for the Bermudan Put option of Table 1 for the stock only strategy ((1(a)) and (1(b))) and the strategy using extra European options (1(c)). Figures (1(d)) and (1(e)) show the P&L with the delta hedging strategy computed explicitly or by Monte-Carlo in the same framework as Figure (1(a)).
First, we comment the results with the stock only portfolio. As we may expect, the rebalancing frequency plays a crucial role. The higher is the frequency, the closer is the estimated dual price to the primal price given by the Longstaff and Schwartz algorithm. We observe on this example the effect of overfitting: for $\bar{N}=10$ and $P=100$, the total number of parameters to estimate is $N\times\bar{N}\times P=10000$, and $Q=100000$ is not sufficient for a suitable estimation, as shown by the difference between $U_{0}^{Q}$ and $\hat{U}_{0}^{Q}$. Increasing $Q$ to $500000$ reduces this gap. With $\bar{N}=20$, $P=50$ and $Q=2\times 10^{6}$, we get a price which is $0.06$ above the Longstaff-Schwartz price, which is satisfactory: the same error magnitude appears also when hedging European options with discrete time strategies.
We now discuss the P&L obtained in Figure 1. We observe the importance of the rebalancing frequency for a stock only portfolio, and get a variance of 2.73 when $\bar{N}=5$ (Figure 1(a)) and $1.05$ when $\bar{N}=10$ (Figure 1(b)). We have also compared the hedging strategy given by our algorithm with the classical delta hedging strategy. In Figure 1(d), we show the P&L distribution obtained with the delta hedging strategy when the delta is computed explicitly with the Cox-Ross-Rubinstein (CRR) approximation of the Black-Scholes model (see Hull (1993) and Pelsser and Vorst (1994)). Compared to Figure 1(a), the distributions are very similar with a slightly sharper hedge for the explicit delta hedging. However, the delta cannot be computed explicitly in general and one has to use approximations method such as the Monte Carlo method proposed by Wang and Caflisch (2010). In Figure 1(e), we have plotted the P&L of the delta hedging obtained with this method when the delta strategy is computed once and for all at the initial time, with a polynomial regression of order 6. The distribution is again quite similar to the one of Figure 1(a), but it is slightly less peaked around zero. These comparisons show that our algorithm produces a sound hedge with respect to classical methods.
When adding the European put option in the hedging portfolio, our algorithm is very close to the primal price with already $\bar{N}=1$ and $P=1$, as expected from the conclusions of Rogers (2002). To be precise, let us consider the following martingale increments between two exercising dates
@@
For the same Bermudan Butterfly option, Table 2 shows the result of our algorithm when using a hedging portfolio using only the asset (${\mathcal{A}}^{1}_{t}=\tilde{S}_{t}$) or using the asset and a European call option (${\mathcal{A}}^{1}_{t}=\tilde{S}_{t}$ and ${\mathcal{A}}^{2}_{t}=e^{-rT}{\mathbb{E}}[(S_{T}-\frac{K_{1}+K_{2}}{2})_{+}|{\mathcal{F}}_{t}]$ in (24)). We see again a clear advantage to include the European option in the hedging strategy and to increase the rebalancing frequency. Thus, our algorithm does not only give a hedging strategy, it can also be used by practitioners to find a good trade-off between the price and the cost of extra monitoring. Besides, we show in Figure 2 that including European options does not only improve the selling price but gives a lower variance of the P&L distribution.